OpenStudy (nicole246):

Write an equation that has a Domain of 1,2,3,4 and a Range of 1,4,9,16

OpenStudy (whpalmer4):

do you know what the question means?

OpenStudy (shadowlegendx):

Input range, get domain Input y, get x f(x) = y f(x) = input \[f(x) = \sqrt x\]

OpenStudy (shadowlegendx):

\[f(1) = \sqrt 1\rightarrow f(1) = 1\]

OpenStudy (shadowlegendx):

\[f(4) = \sqrt 4 \rightarrow f(4) = 2\]

OpenStudy (shadowlegendx):

\[f(9) = \sqrt 9 \rightarrow f(9) = 3\]

OpenStudy (shadowlegendx):

\[f(16) = \sqrt 16 \rightarrow f(16) = 4\]

OpenStudy (whpalmer4):

you have them switched, shadow legend

OpenStudy (whpalmer4):

domain is the set of X values range is the set of Y values

OpenStudy (shadowlegendx):

x^2 then

OpenStudy (shadowlegendx):

\[f(x) = x^2\]

OpenStudy (whpalmer4):

the function here to convert 1,2,3,4 to 1,4,9,16 is what?

OpenStudy (shadowlegendx):

To be fair, I was typing this solution whilst talking about food, whilst being hungry e-e

OpenStudy (unklerhaukus):

squares!

OpenStudy (mathmale):

What U. R. is trying to say is this: Take each element of the domain, and then, one by one, square that element. What output values do you obtain? If you do this correctly, you will have produced the "range" of this function.